OpenStudy (anonymous):

Identify all of the unknowns in the problem. There are 5 more girls than boys in a class. There are 27 students in the class in all. How many boys are there in the class? (Points : 1) the number of students in the class the number of boys in the class the number of girls in the class the number of boys and the number of girls in the class

OpenStudy (anonymous):

D

OpenStudy (anonymous):

11 boys, 16 girls, 27 students, 27 boys and girls

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

Zach is 4 years older than twice his sister Maya’s age. If m is Maya’s age, write an expression that represents Zach’s age. (Points : 1) 2m + 4 m + 4 2m 4m + 2

OpenStudy (anonymous):

2m +4

OpenStudy (anonymous):

thank you your really smart at this :)

OpenStudy (anonymous):

Melanie is 2 years older than twice her brother Bill’s age. If Bill is 5 years old, how old is Melanie? (Points : 1) 4 10 12 14

OpenStudy (anonymous):

12

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Write an equation to model the given situation. The area of a rectangle is 60 square inches. The length of the rectangle is 7 inches longer than the width. (Points : 1) w + 7 = 60 7w = 60 w(w + 7) = 60 w + 7w = 60

OpenStudy (anonymous):

C

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Solve the given problem. Nia is three years older than half Jordan’s age. If Jordan is eight years old, how old is Nia? (Points : 1) 19 11 4 7

OpenStudy (anonymous):

the question appears incomplete to me.

OpenStudy (anonymous):

7 i think

OpenStudy (anonymous):

7

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

i wish I could give you both medals you guys are so smart

OpenStudy (anonymous):

Just hit best response on one of our comments. You can also click on our name and become a fan

OpenStudy (anonymous):

i will give cody the medal for you ok.

OpenStudy (anonymous):

aw thanks

OpenStudy (anonymous):

I gave you one @kevindragonlord

OpenStudy (anonymous):

do you guys go to GCA or to any school belonging to K12?

OpenStudy (anonymous):

I go to TRECA its a homeschool